OpenStudy (anonymous):

is 5/x^-4 already simplified?

OpenStudy (e.mccormick):

\[\frac{5}{x^{-4}} \]Not simplified. \[5x^{4}\]Simplified.

OpenStudy (anonymous):

How do you simplify it?

OpenStudy (e.mccormick):

Don't know why you deleted the other answer... you had it.

OpenStudy (e.mccormick):

The rules of exponents: \[\frac{1}{a^{-n}}=a^n\]

OpenStudy (e.mccormick):

Also: \[a^{-n}=\frac{1}{a^{n}}\]

OpenStudy (anonymous):

Ok. Thanks

OpenStudy (e.mccormick):

If you remember that a negative exponent causes something to jump to the other side of a quotent, you can get pretty much any of these right. Don't have too much fun!